# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

### v1

@@ 3.3 Fixed-Lag Strategy for the Online Smoother
The fixed-lag smoother is the simplest method for conserving computational and storage resources, as it assumes a uniformly predetermined impact region for each new observation. Setting an a priori impact region incurs no additional computational cost. However, this approach has significant drawbacks for most turbulent and complex dynamical systems: it may assign an unnecessarily long fixed lag, wasting storage with diminishing returns in state estimation, or a short fixed lag that introduces substantial biases in the estimated state. Nevertheless, it serves as a building block of the adaptive lag smoother.
@@
The fixed-lag methodology is defined as follows: we assume a predetermined impact region size for each new observation, denoted as lag $L\in\{0,\ldots,n\}$ in time steps (edge cases $L=0$ and $L=n$ are discussed later). All subsequent measurements can be expressed in simulation time units by multiplying by $\Delta t$. We also assume that $L$ is uniform across $n$, the number of observations. The fixed-lag online smoother Gaussian statistics are then given by the following formulas:
$\boldsymbol{\mu}_{\text{s}}^{j,n}$ :=$:=\begin{cases}\boldsymbol{\mu}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L\\ \boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big{(}\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big{)},&\text{ for }n-L\leq j\leq n-1,\end{cases}$ (3.23)
$\mathbf{R}_{\text{s}}^{j,n}$ :=$:=\begin{cases}\mathbf{R}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L\\ \mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big{(}\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\big{)}(\mathbf{D}^{j,n-2})^{\dagger},&\text{ for }n-L\leq j\leq n-1,\end{cases}$ (3.24)
@@
with the details for the online discrete smoother already outlined in Algorithm 1 of Section 3.1. For $j=n$, as already discussed, we always have $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$. As such, we have for $\forall n\in\mathbb{N}$ and $j=0,\ldots,n-1$, that $\boldsymbol{\mu}_{\text{s}}^{j,n}=\boldsymbol{\mu}_{\text{s}}^{j,n}(L)$ and $\mathbf{R}_{\text{s}}^{j,n}=\mathbf{R}_{\text{s}}^{j,n}(L)$, where this explicitly denotes the dependence on the predetermined fixed lag value of $L$.
@@
We mention here that the edge cases $L=0$ and $L=n$ actually recover the optimal filter and (offline) smoother state estimations, respectively, under the consideration that we consider the $n$ observations to define a complete time series for the observable variables. It is not hard to see why this is the case. Observe that when $L=0$, then we only work with the first branch of the definitions for the fixed-lag smoother Gaussian statistics in (3.23) and (3.24), and as such with each new observation we just carry over the estimated state from the corresponding time instant $t_{j}$ we did at the previous iteration, i.e., the one from the last observation. As such, since $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$, for every $n\in\mathbb{N}_{0}$, then it is immediate that when $L=0$, then with each new observation we just carry-on with the forward-pass and calculate the optimal filter Gaussian statistics at the new observation and then carry over all the previous state estimates. But since at each point we are calculating the filter state, then we are essentially just doing a forward-pass as the observations come in, and as such we only recover the filter posterior Gaussian statistics. On the other hand, when $L=n$, then we only work with the second branch of the definitions for the fixed-lag smoother Gaussian statistics in (3.23) and (3.24), which means that with each new observation we do the forward-pass and obtain the new filter estimate at the end-point ($j=n$), but using this new state estimation we do a full backward-pass update and obtain the smoother estimates at each and every time instant, when now considering this new observation. As such, we are essentially doing a full forward- and backward-pass with the arrival of each observation, which means we are recovering with each new observational data the offline smoother posterior Gaussian statistics, fully, for this complete observational period.
@@ 3.4 Online Smoother with an Adaptive Lag Determined Using Information Theory
3.4.1 Information Theory Fundamentals: Relative Entropy and the Signal-Dispersion Decomposition
Due to the turbulent nature, new observational data influence the estimated states through the online smoother update only within a finite time interval, with this impact exhibiting exponential decay over time. Therefore, it is natural to quantify the information gain from applying the optimal online forward-in-time discrete smoother (posterior) with different length of lags compared to simply carrying forward the previous state estimate (prior). By doing so, the optimal lag can be discovered. We utilize empirical information theory to develop a measure that quantifies this information gain, reflecting the additional information in the posterior distribution beyond the prior [87, 88, 89, 90, 91].
In the information-theoretic framework, a natural way to assess the information gain in the posterior when compared to the prior is through their respective PDFs by using a certain type of distance function that measures the difference between them. A general class of such functions is known as $f$-divergences [92, 93, 94].
While the general framework of $f$-divergences is rather versatile, there is a certain choice of $f$ which takes great advantage of the form of the CGNS framework, and especially of the conditional Gaussianity of the posterior distributions. Specifically, by choosing $f(z)=z\log z$, then we recover the so called Kullback-Leibler divergence or relative entropy [95, 96, 97]. In this case, we denote the divergence with $\mathcal{P}(\boldsymbol{\cdot},\boldsymbol{\cdot})$, and is simply given as [98, 99, 100]
@@ 3.4.1 Information Theory Fundamentals: Relative Entropy and the Signal-Dispersion Decomposition
where the integration with respect to the state vector is to be understood in the Lebesgue sense and is essentially taken over the common support of the two densities. For our framework, we read $\mathcal{P}(p,q)$ as the information gain of $p$ from $q$, as in the information that has been gained by using the true density $p$ in lieu of the approximation $q$. The relative entropy enjoys the lucrative properties of nonnegativity and of invariancy under general nonlinear changes of state variables [101]. These properties, among others, allow the relative entropy to be widely utilized in assessing the information gain and quantifying model error, model sensitivity, predictability, and statistical response for stochastic CTNDSs [102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112].
The relative entropy ensures that the extreme events related to tail probabilities are not underestimated. The ratio of the two PDFs within the logarithm quantifies the gap in the tail probability, allowing for an unbiased characterization of statistical differences. Many CTNDSs modeled through the CGNS framework generate intermittent, rare, and extreme events. Thus, using a logarithmic generator function effectively resolves tail probability events, aiding in quantifying information gain for the adaptive lag criterion, uncertainty reduction of posterior solutions, and causal inference. In addition, relative entropy benefits from this logarithmic structure, providing a simple, closed-form formula when both distributions are Gaussian. This is ideal for the online smoother update, where both the prior and posterior distributions are Gaussian, and differ only in their statistics. Consequently, this facilitates the efficient calculation of adaptive lag, reducing computational costs and storage requirements. Specifically, when both $\mathbb{P}\overset{{\,\rm d}}{\sim}\mathcal{N}_{N}(\boldsymbol{\mu}_{\text{p}},\mathbf{R}_{\text{p}})$ and $\mathbb{Q}\overset{{\,\rm d}}{\sim}\mathcal{N}_{N}(\boldsymbol{\mu}_{\text{q}},\mathbf{R}_{\text{q}})$, where $N=\mathrm{dim}(\mathbf{u})$ is the dimension of the phase space, then the relative entropy adopts an explicit and simple formula [2, 105],
@@
where $\mathrm{tr}$ and $\mathrm{det}$ are the trace and determinant of a matrix, respectively. The first quadratic form term in (3.26) is called the signal and measures the information gain in the mean weighted by the model or approximation covariance, whereas the second term is called the dispersion and involves only the covariance ratio $\mathbf{R}_{\text{p}}\mathbf{R}_{\text{q}}^{-1}$. This is why the expression in (3.26) is known as the signal-dispersion decomposition of the relative entropy for Gaussian variables.
@@ 3.4.2 Development of the Adaptive-Lag Strategy for the Online Smoother
Here we outline the information-theoretic approach for defining the adaptive lag at the acquirement of a new observation. Let us assume that we have currently observed the $n$-th measurement, $\mathbf{x}^{n}$, which is used to update all the existing online smoother posterior Gaussian statistics at time instants $10\leq j\leq n-$ 1, and then provide a new state estimate at $j=n$. As shown in Theorem 3.1, that update happens via (3.11)–(3.12). We would like to define an adaptive lag at the $n$-observation, $L_{n}\in\{1,\ldots,n-1\}$, such that only “the most informative or impactful” of these updates are actually carried out, in other words,
$\boldsymbol{\mu}_{\text{s}}^{j,n}\equiv\boldsymbol{\mu}_{\text{s}}^{j,n}(L_{n})$ :=$:=\begin{cases}\boldsymbol{\mu}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L_{n}\\ \boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big{(}\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big{)},&\text{ for }n-L_{n}\leq j\leq n-1,\end{cases}$ (3.27)
$\mathbf{R}_{\text{s}}^{j,n}\equiv\mathbf{R}_{\text{s}}^{j,n}(L_{n})$ :=$:=\begin{cases}\mathbf{R}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L_{n}\\ \mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big{(}\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\big{)}(\mathbf{D}^{j,n-2})^{\dagger},&\text{ for }n-L_{n}\leq j\leq n-1,\end{cases}$ (3.28)
with the details for the online discrete smoother already outlined in Algorithm 1 of Section 3.1. For $j=n$, as already discussed, we always have $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$.
To define the adaptive lag $L_{n}$ using the information theory principles from Section 3.4.1, we quantify the information gain from updating the state estimate at the $j$-th time step with the $n$-th observation, compared to using the estimate from the previous $(n-1)$-st observation. In the CGNS framework, two possible Gaussian distributions arise for the online smoother state estimation. The first, $p^{j,n}_{\text{updated}}$, is the optimal posterior distribution obtained through the online forward-in-time smoother, discussed in Section 3.1. The second, $p^{j,n}_{\text{lagged}}$, is the prior distribution, where no update is made and the previous state estimate at time $t_{j}$ is carried over instead. The posterior or the “updated” distribution has Gaussian statistics given by
$\boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right)\quad\text{and}\quad\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-2})^{\dagger},$
@@
i.e., no update (or innovation, under the Kalman gain terminology) is being made or procured due to the newly acquired observation. This is precisely described by the first and second branches of (3.27)–(3.28), respectively.
With the posterior or “updated” conditional Gaussian distribution denoting the true distribution, while the prior or “lagged” conditional Gaussian distribution denotes the approximation due to the adaptive lag (by not carrying out the online smoother update), we can then naturally take advantage of the signal-dispersion decomposition formula for the relative entropy given in (3.26). As such, we have that the information gain or uncertainty reduction incurred by carrying out the update through the online smoother at time $t_{j}$ by using the newly obtained $n$-th observation, is equal to
$\begin{split}\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)=&\ \frac{1}{2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right)^{\dagger}(\mathbf{D}^{j,n-2})^{\dagger}(\mathbf{R}_{\text{s}}^{j,n-1})^{-1}\mathbf{D}^{j,n-2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right)\\ &+\frac{1}{2}\left(\mathrm{tr}(\mathbf{Q}^{j,n})-l-\mathrm{ln}(\mathrm{det}(\mathbf{Q}^{j,n}))\right),\end{split}$ (3.29)
where $\mathbf{Q}^{j,n}$ is defined as a covariance ratio matrix,
$\begin{split}\mathbf{Q}^{j,n}&:=\left(\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-2})^{\dagger}\right)(\mathbf{R}_{\text{s}}^{j,n-1})^{-1}\\ &=\mathbf{I}_{l\times l}+\left(\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-2})^{\dagger}\right)(\mathbf{R}_{\text{s}}^{j,n-1})^{-1}.\end{split}$ (3.30)
@@
[…]
@@
Notice that for larger values of $\delta$ we allow for more relaxed or smaller lags, while for smaller tolerance values we push the adaptive lag values towards their predefined upper bound, $b$. Per our methodology, and as noted in (3.31), we have that $\forall n\in\mathbb{N}$, $L_{n}\in\{0,1,\ldots,b\}$.
@@
where again if no such maximiser exists we simply use (3.32). For most cases (cf. the numerical experiments in Sections 4.1 and 4.2 where $\sigma^{j,n}$ and $\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)$ behave as $C^{n}_{1}e^{C^{n}_{2}j}$ for $R_{n}\leq j\leq n-1$ and $C^{n}_{1},C^{n}_{2}>0$) leads to the same results after a slight modification to the tolerance parameter being used (and for appropriate values of $w$ in (3.31)). In such settings, if we use $\delta=O(10^{-d})$ with the sequence of local standard deviations, for $d\in\mathbb{N}$, then we can get similar results when using the original sequence of relative entropies by using $\delta=O(10^{-d+s/2})$ where $s\in\mathbb{N}$ is defined through the time step as $\Delta t=O(10^{-s})$, since the former is approximately the first derivative of the latter with respect to the time step, as already mentioned before. The simplicity and flexibility of the adaptive lag framework in (3.31) or (3.33) further allows for the inclusion of a penalty for large lag values. This can be achieved by modifying the criterion (e.g., by using a monotonic polynomial of time as the penalization) and then choosing the largest lag for which this penalized information gain (or its penalized local standard deviation) remains small under the tolerance $\delta$, while prioritizing smaller lags.

### v2

@@ 3.3 Fixed-Lag Strategy for the Online Smoother
The fixed-lag smoother is the simplest method for conserving computational and storage resources, as it assumes a uniformly predetermined impact region for each new observation. Setting an a-priori impact region incurs no additional computational cost. However, this approach has significant drawbacks for most turbulent and complex dynamical systems: it may assign an unnecessarily long fixed lag, wasting storage with diminishing returns in state estimation, or a short fixed lag that introduces substantial biases in the estimated state. Nevertheless, it serves as a building block for the adaptive-lag online smoother.
@@
The fixed-lag methodology is defined as follows: we assume a predetermined impact region size for each new observation, denoted as lag $L\in\{1,\ldots,n-1\}$ in time steps (the edge cases $L=0$ and $L=n$ are discussed later). All subsequent measurements can be expressed in simulation time units by multiplying by $\Delta t$. We also assume that $L$ is uniform across $n$, the number of observations. The fixed-lag online smoother Gaussian statistics are then given by the following formulae:
$\boldsymbol{\mu}_{\text{s}}^{j,n}\equiv\boldsymbol{\mu}_{\text{s}}^{j,n}(L)$ $:=\begin{cases}\boldsymbol{\mu}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L\\ \boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big),&\text{ for }n-L\leq j\leq n-1,\end{cases}$ (3.22)
$\mathbf{R}_{\text{s}}^{j,n}\equiv\mathbf{R}_{\text{s}}^{j,n}(L)$ $:=\begin{cases}\mathbf{R}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L\\ \mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\big)(\mathbf{D}^{j,n-2})^{\dagger},&\text{ for }n-L\leq j\leq n-1,\end{cases}$ (3.23)
@@
with the details of the online discrete smoother being outlined in Algorithm 1 of Section 3.1. For $j=n$, as already discussed, we always have $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$. As such, we have for $\forall n\in\mathbb{N}$ and $j=0,\ldots,n-1$, that $\boldsymbol{\mu}_{\text{s}}^{j,n}=\boldsymbol{\mu}_{\text{s}}^{j,n}(L)$ and $\mathbf{R}_{\text{s}}^{j,n}=\mathbf{R}_{\text{s}}^{j,n}(L)$, where this explicitly denotes the dependence on the predetermined fixed lag value of $L$.
@@
We mention here that the edge cases of $L=0$ and $L=n$ actually recover the optimal filter and (offline) smoother state estimations, respectively, under the consideration that the current $n$ observations define a complete time series for the observable variables. It is not hard to see why this is the case. Observe that when $L=0$, then we only work with the first branch of the definitions for the fixed-lag smoother Gaussian statistics in (3.22) and (3.23), and as such with each new observation we just carry over the estimated state from the corresponding time instant $t_{j}$ at the previous iteration, i.e., from the last observation. As such, since $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$, for every $n\in\mathbb{N}_{0}$, then it is immediate that when $L=0$, with each new observation we just carry-on with the forward-pass and calculate the optimal filter Gaussian statistics at the new observation and then carry over all the previous state estimates. But since at each point we are calculating just the filter state, then we are essentially doing a forward-pass as the observations come in, and as such we only recover the filter Gaussian statistics. On the other hand, when $L=n$, then we only work with the second branch of the definitions for the fixed-lag smoother Gaussian statistics in (3.22) and (3.23), which means that with each new observation we continue the forward-pass and obtain the new filter estimate at the end-point ($j=n$), but using this new state estimation we do a full backward-pass update and obtain the smoother estimates at each and every time instant. As such, we are essentially doing a full forward- and backward-pass with the arrival of each observation, which means we are recovering with each new observational data the offline smoother posterior Gaussian statistics, fully, for this complete observational period.
@@ 3.4 Online Smoother with an Adaptive Lag Determined Using Information Theory
3.4.1 Information Theory Fundamentals: Relative Entropy and the Signal–Dispersion Decomposition
Due to the turbulent dynamics, new observational data influence the estimated states through the online smoother update only within a finite time interval, with this impact exhibiting exponential decay over time. Therefore, it is natural to quantify the information gain from applying the optimal online forward-in-time discrete smoother (posterior), with varying lengths of lags, compared to simply carrying forward the previous state estimate (prior). By doing so, the optimal lag can be discovered. We utilize empirical information theory to develop a measure that quantifies this information gain [70], reflecting the additional information in the posterior distribution beyond the prior [71, 72, 73]. In the information-theoretic framework, this is naturally achieved through their respective PDFs, by using a certain type of distance function that measures the statistical discrepancy between them. A general class of such functions is known as $f$-divergences [74, 75].
While the general framework of $f$-divergences is rather versatile, there is a certain choice of $f$ which takes great advantage of the form of the CGNS framework, and especially of the conditional Gaussianity of the posterior distributions. Specifically, by choosing $f(z)=z\log z$, then we recover the so called Kullback–Leibler divergence or relative entropy [76, 77]. In this case, we denote the divergence with $\mathcal{P}$, and is simply given as [78, 79]
@@ 3.4.1 Information Theory Fundamentals: Relative Entropy and the Signal–Dispersion Decomposition
where the integration with respect to the state vector is to be understood in the Lebesgue sense [^3] We implicitly assume in (3.24) that $p$ and $q$ are the PDFs of two probability distributions, $\mathbb{P}$ and $\mathbb{Q}$, respectively, with $\mathbb{P}\ll\mathbb{Q}$ (i.e., $\mathbb{P}$ is dominated by $\mathbb{Q}$), so the integration is taken over the support of $q$, where both $\mathbb{P}$ and $\mathbb{Q}$ are dominated by the Lebesgue measure.. For our framework, we read $\mathcal{P}(p,q)$ as the information gain of $p$ from $q$, as in the information that has been gained by using the true density $p$ in lieu of the approximation $q$. The relative entropy enjoys the lucrative properties of non-negativity and of invariance under general nonlinear changes of state variables [80]. These properties, among others, allow the relative entropy to be widely utilized in assessing the information gain and quantifying model error, model sensitivity, predictability, and statistical response for stochastic CTNDSs [81, 82, 83, 84, 85, 86, 87].
The relative entropy ensures that the extreme events related to tail probabilities are not underestimated. The ratio of the two PDFs within the logarithm quantifies the gap in the tail probability, allowing for an unbiased characterization of statistical differences. Many CTNDSs modeled through the CGNS framework generate intermittent, rare, and extreme events. Thus, using a logarithmic generator function effectively resolves tail probability events, aiding in quantifying information gain for the adaptive lag criterion, uncertainty reduction of posterior solutions, and causal inference. In addition, relative entropy benefits from this logarithmic structure, providing a simple, closed-form formula when both distributions are Gaussian. This is ideal for the online smoother update, where both the prior and posterior distributions are Gaussian, and differ only in their statistics. Consequently, this facilitates the efficient calculation of the adaptive lag, thus further reducing computational costs and storage requirements. Specifically, when both $\mathbb{P}\overset{{\,\rm d}}{\sim}\mathcal{N}_{N}\big(\boldsymbol{\mu}_{\text{p}},\mathbf{R}_{\text{p}}\big)$ and $\mathbb{Q}\overset{{\,\rm d}}{\sim}\mathcal{N}_{N}\big(\boldsymbol{\mu}_{\text{q}},\mathbf{R}_{\text{q}}\big)$, where $N=\mathrm{dim}(\mathbf{u})$ is the dimension of the phase space, then the relative entropy adopts an explicit and simple formula [2, 82]:
@@
where $\mathrm{tr}(\cdot)$ and $\mathrm{det}(\cdot)$ are the trace and determinant of a matrix, respectively. The first quadratic form term in (3.25) is called the signal and measures the information gain in the mean weighted by the model or approximation covariance, whereas the second term is called the dispersion and involves only the covariance ratio $\mathbf{R}_{\text{p}}\mathbf{R}_{\text{q}}^{-1}$. This is why the expression in (3.25) is known as the signal–dispersion decomposition of the relative entropy for Gaussian variables.
@@ 3.4.2 Development of the Adaptive-Lag Strategy for the Online Smoother
Here we outline the information-theoretic approach for defining the adaptive lag at the acquirement of a new observation. Let us assume that we have currently observed the $n$-th measurement, $\mathbf{x}^{n}$, which is used to update all the existing online smoother Gaussian statistics at time instants $0\leq j\leq n-1$, and then provide a new state estimate at $j=n$. As shown in Theorem 3.1, that update happens via (3.10)–(3.11). We would like to define an adaptive lag at the $n$-observation, $L_{n}\in\{1,\ldots,n-1\}$, such that only “the most informative or impactful” of these updates are actually carried out, in other words,
$\boldsymbol{\mu}_{\text{s}}^{j,n}\equiv\boldsymbol{\mu}_{\text{s}}^{j,n}(L_{n})$ $:=\begin{cases}\boldsymbol{\mu}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L_{n}\\ \boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big),&\text{ for }n-L_{n}\leq j\leq n-1,\end{cases}$ (3.26)
$\mathbf{R}_{\text{s}}^{j,n}\equiv\mathbf{R}_{\text{s}}^{j,n}(L_{n})$ $:=\begin{cases}\mathbf{R}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L_{n}\\ \mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\big)(\mathbf{D}^{j,n-2})^{\dagger},&\text{ for }n-L_{n}\leq j\leq n-1,\end{cases}$ (3.27)
with the details of the online discrete smoother being outlined in Algorithm 1 of Section 3.1. For $j=n$, as already discussed, we always have $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$.
To define the adaptive lag $L_{n}$ using the information theory principles from Section 3.4.1, we quantify the information gain from updating the state estimate at the $j$-th time step with the $n$-th observation, compared to using the estimate from the previous (n1)(n-1)-st observation. In the CGNS framework, two possible Gaussian distributions arise for the online smoother state estimation. The first, $p^{j,n}_{\text{updated}}$, is the optimal posterior distribution obtained through the online forward-in-time smoother, discussed in Section 3.1. The second, $p^{j,n}_{\text{lagged}}$, is the prior distribution, where no update is made and the previous state estimate at time $t_{j}$ is carried over instead. The posterior or the “updated” distribution has Gaussian statistics given by
$\boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big)\quad\text{and}\quad\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\big)(\mathbf{D}^{j,n-2})^{\dagger},$
@@
i.e., no update (or innovation, under the Kalman gain terminology) is being made or procured due to the newly acquired observation. This is precisely described by the first branch in (3.26) and (3.27).
With the posterior or “updated” conditional Gaussian distribution denoting the true distribution, while the prior or “lagged” conditional Gaussian distribution denotes the approximation due to the adaptive lag (by not carrying out the online smoother update), we can then naturally take advantage of the signal–dispersion decomposition formula for the relative entropy given in (3.25). As such, we have that the information gain or uncertainty reduction incurred by carrying out the update through the online smoother at time $t_{j}$ by using the newly obtained $n$-th observation, is equal to
$\begin{split}\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)=&\ \frac{1}{2}\big(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big)^{\dagger}(\mathbf{D}^{j,n-2})^{\dagger}(\mathbf{R}_{\text{s}}^{j,n-1})^{-1}\mathbf{D}^{j,n-2}\big(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big)\\ &+\frac{1}{2}\big(\mathrm{tr}(\mathbf{Q}^{j,n})-l-\mathrm{ln}(\mathrm{det}(\mathbf{Q}^{j,n}))\big),\end{split}$ (3.28)
where $\mathbf{Q}^{j,n}$ is defined as a covariance ratio matrix
$\begin{split}\mathbf{Q}^{j,n}&:=\big(\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-2})^{\dagger}\big)(\mathbf{R}_{\text{s}}^{j,n-1})^{-1}\\ &=\mathbf{I}_{l\times l}+\big(\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-2})^{\dagger}\big)(\mathbf{R}_{\text{s}}^{j,n-1})^{-1}.\end{split}$ (3.29)
@@
values. Two possible approaches to define the adaptive lag, both yielding similar results after adjusting hyperparameters, are: we can either use the sequence of relative entropies $\left\{\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)\right\}_{R_{n}\leq j\leq n-1}$, or apply the local standard deviation (LSDev) $\{\sigma^{j,n}\}_{R_{n}\leq j\leq n-1}$ to the original entropy sequence and use that instead. The LSDev of the entropy sequence serves as a proxy for the first derivative of the information gain (when estimated via a finite difference scheme), efficiently identifying regions of stagnation or rapid change in it. Specifically $\sigma^{j,n}$, for a moving window span size of $w\equiv 1(\mathrm{mod}\ 2)$, is the standard deviation of the neighborhood centered at the corresponding input value (or pixel in image filtering) over $j$, extending (w1)/2(w-1)/2 positions to the left and to the right, for each $n$. This can potentially allow for more effective adaptive lag selection. Both methods behave similarly in most numerical experiments, including all case studies in Section 4, especially when the information gain sequence behaves exponentially with respect to $j$ for each $n$. In such cases, both the relative entropy and its LSDev exhibit the same growth, making the approaches roughly equivalent. For simplicity, in what follows, we present the adaptive lag procedure using the LSDev. However, the original relative entropy sequence of information gains could also be used directly. Naturally, when using $\{\sigma^{j,n}\}_{R_{n}\leq j\leq n-1}$, we additionally have to consider the dependence of the procedure on the window size being used to define the moving LSDev. Taking this into consideration, for a moving window span size of $w\in\mathbb{N}_{\geq 3}$, with $w\equiv 1(\mathrm{mod}\ 2)$, which gives $\{\sigma^{j,n}\}_{R_{n}\leq j\leq n-1}=\{\sigma^{j,n}(w)\}_{R_{n}\leq j\leq n-1}$, and a tolerance parameter $\delta\in(0,1)$, the definition of the adaptive lag at the $n$-th observation in this framework is given as
@@
(Throughout all numerical case studies in this work, we set $w=7$ which is MATLAB’s default value for stdfilt().) Notice that for larger values of $\delta$ we allow for more relaxed or smaller lags, while for smaller tolerance values we push the adaptive lag values towards their predefined upper bound, $b$. Per our methodology, and as noted in (3.30), we have that $\forall n\in\mathbb{N}$, $L_{n}\in\{0,1,\ldots,b\}$.
@@
where again if no such maximiser exists we simply use (3.31). For most cases this approach leads to the same results after a slight modification to the tolerance parameter being used (and for appropriate values of $w$ in (3.30); see the numerical results in Figure 4.5 of Section 4.2, where both $\sigma^{j,n}$ and $\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)$ behave as $C^{n}_{1}e^{C^{n}_{2}j}$ for $R_{n}\leq j\leq n-1$ and $C^{n}_{1},C^{n}_{2}>0$). In such settings, if we use $\delta=O(10^{-d})$ with the sequence of local standard deviations, for $d\in\mathbb{N}$, then we can get similar results when using the original sequence of relative entropies by using $\delta=O(10^{-d+s/2})$ where $s\in\mathbb{N}$ is defined through the time step as $\Delta t=O(10^{-s})$, since the former is approximately the first derivative of the latter with respect to the time step, as already mentioned. Furthermore, the simplicity and flexibility of the adaptive lag framework in (3.30) or (3.32) further allows for the inclusion of a penalty for large lag values. This is achieved by modifying the criterion (e.g., by using a monotonic polynomial of time as the penalization) and then choosing the largest lag for which this penalized information gain (or its penalized local standard deviation) remains small under the tolerance $\delta$, while prioritizing smaller lags.
[…]